OpenStudy (anonymous):

What is the volume of the cube with side length 15 in. ? choices: 90 cu in. 1,350 cu in. 3,375 cu in. 45 cu in.

OpenStudy (anonymous):

it was already answered>.....@HiImOlivia v=s^3

OpenStudy (anonymous):

try the formula

OpenStudy (anonymous):

multiply the number by itself three times \[15^3=15\times 15\times 15\]

OpenStudy (anonymous):

thanks @satellite73

OpenStudy (anonymous):

yw